The diameter of the round brass rod is 30.05 mm and the steel ring the inner diameter is 30.00 mm at 20 ° C. At what temperature the steel ring must be heated to fit the brass rod inside?
1
Expert's answer
2019-03-18T11:38:54-0400

Solution:

Change in length of the ring with temperature increasing:


"\Delta{L}=L_0\alpha\Delta{T}"

Then change in temperature:


"\Delta{T}=\frac{\Delta{L}} {L_0\alpha}=\frac{L-L_0} {L_0\alpha}"

Length of the ring after temperature increasing:

"L=2\pi R=\frac{2\pi D} 2=\pi D"

Analogously we can find the length of the ring before temperature increasing:

"L_0=\pi D_0"

Therefore:

"\Delta{T}=\frac{\pi D-\pi D_0} {\pi D_0\alpha}=\frac{D-D_0} {D_0\alpha}"

Coefficient of linear expansion for steel:


"\alpha=1.2\times 10^{-5} \: (1/^\circ C)"

Let's calculate the change in temperature:


"\Delta{T}=\frac{30.05-30.00} {30.00\times 1.2\times 10^{-5}}=139\:(^\circ\:C)"

Therefore  the temperature the steel ring must be heated to fit the brass rod inside:

"T_2=T_1+\Delta{T}=20^\circ\:C+139^\circ\:C=159\:^\circ\:C"

Answer: "T_2=159\:^\circ\:C".